efer to the following information about the Finishing Department in the Gallagher Factory for the month of June. Gallagher Factory uses the FIFO method of inventory costing.

Beginning Work in Process inventory:

Physical units…………………………………………... 5,000 units
% complete for materials 70%
% complete for conversion costs 25%
Materials cost from May $7,350
Conversion costs from May $3,125

Product started:

Physical units 44,000 units

Ending Work in Process inventory:

Physical units 4,000 units
% complete for materials 40%
% complete for conversion costs 10%

Manufacturing costs for June:

Materials $96,975
Conversion costs $79,470

Compute costs in ending inventory for June.
Answer #1
Costs in ending inventory for June 4320
Workings:
Materials Conversion costs
To complete Beginning Work in Process inventory 1500 3750
Units started and completed 40000 40000
Ending Work in Process inventory 1600 400
Equivalent units 43100 44150
Manufacturing costs for June 96975 79470
/ Equivalent units 43100 44150
Cost per Equivalent unit 2.25 1.80
Materials Conversion costs Total
Ending Work in Process inventory 1600 400
X Cost per Equivalent unit 2.25 1.80
Cost of Ending Work in Process inventory 3600 720 4320
Materials Conversion costs
To complete Beginning Work in Process inventory =5000*(1-70%) =5000*(1-25%)
Units started and completed =44000-4000 =44000-4000
Ending Work in Process inventory =4000*40% =4000*10%